# LINSCAN - A Linearity Based Clustering Algorithm

### v1

@@ 3.3 Local Behavior of $D$
The proof relies on a significant number of inequalities and is provided in appendix B. The proof proceeds by separating the first two terms of $D(P,K)$ from the last two and showing that each pair individually satisfies the triangle inequality with small additive errors.
Importantly, this shows that for small values of $\epsilon$, $D$ behaves approximately like a metric, which allows us to bound the diameter of any cluster in terms of $\epsilon$ and the number of steps between points in the cluster. This ensures that points whose local neighborhoods are nearly orthogonal are not clustered together.
Compare this to the best results proven previously for the approximate triangle inequality of the unmodified KL-Divergence between Gaussians in [12], which was of exponential order.
@@ 4 Numerical Results
Experiments with synthetic data sets revealed that some clusters identified by LINSCAN may not appear as sufficiently ”linear” upon visual inspection (e.g., due to high scatter of data points). Therefore we introduce an additional quality check whereby we compute the covariance matrix of each cluster. In the case of $\mathbb{R}^{2}$ we set a minimal threshold on the ratio of the minimum eigenvalue to the maximum eigenvalue of the covariance matrix and discard clusters which do not meet this threshold.
@@ 4.1 Example Datasets
Figure 3(a) shows the results of applying LINSCAN to the same data as in Figure 1(a), Figure 3(b) shows the clusters with the noise points removed, and Figure 3(c) shows the results of removing clusters with spectral ratio greater than $\frac{1}{2}$. Note the separation of clusters in the bottom left and top left corners in comparison to the results from DBSCAN.
Figure 4 shows the results of applying LINSCAN to real data representing earthquake epicenters in Southern California [13]. Not only does LINSCAN identify QLCs and removes the ”diffuse” background seismicity, but it is also able to identify the clusters at multiple distinct scales simultaneously.
Figure 2 shows the example of synthetic data with two QLCs intersecting at a high angle. Note that LINSCAN is the only algorithm that is able to distinguish between the two lineated sets of data points.
@@
Figure 3: Synthetic Data
(a) LINSCAN Results
(b) LINSCAN Results with Noise Removed
(c) LINSCAN Results with Isotropic Clusters Removed
@@
(b) LINSCAN results
(c) LINSCAN Results with Noise Removed
(d) LINSCAN Results with Isotropic Clusters Removed
@@ 4.2 Measuring Performance
To quantitatively evaluate the algorithm performance, we conducted several tests on synthetic, labeled data. We generated 10 linear clusters, 5 isotropic clusters, and 10 pairs of linear clusters intersecting at angles in the range [.$1\pi,.9\pi][ .$ and separated them from one another in space. An example is given in Figure 5(a).
To score the performance, we use the Adjusted Rand Index from [14]:
Definition 1 (Rand Index)
Let $X=\left\{x_{1},...,x_{n}\right\}$ be a set of elements and consider two partitions $\mathcal{C}=\left\{C_{1},...,C_{m}\right\}$ and $\mathcal{C}^{\prime}=\left\{C^{\prime}_{1},...,C^{\prime}_{n}\right\}$, i.e. $C_{i}\subseteq X$ and $C_{i}^{\prime}\subseteq X$ for all $i$ and
$X=\bigcup_{i=1}^{m}C_{i}=\bigcup_{i=1}^{n}C^{\prime}_{i}$
with
$C_{i}\cap C_{j}=C_{i}^{\prime}\cap C_{j}^{\prime}=\emptyset$
for all $i\neq j$. Let
$a:=\#\left\{(x,y)\in X\times Y:x\neq y,\exists i,j\text{ s.t. }x,y\in C_{i},x,y\in C_{j}^{\prime}\right\}$
and
$b:=\#\left\{(x,y)\in X\times Y:x\neq y,\exists i,j,k,l\text{ s.t. }i\neq j,k\neq l,x\in C_{i},x\in C_{k}^{\prime},y\in C_{j},y\in C_{l}^{\prime}\right\}$
In words, $a$ is the number of pairs of elements of $X$ such that both elements are in the same cluster in $\mathcal{C}$ and $\mathcal{C}^{\prime}$ and $b$ is the number of pairs of elements of $X$ such that both elements are in different clusters in both $\mathcal{C}$ and $\mathcal{C}^{\prime}$. Then, the Rand Index is given by
$R(X,\mathcal{C},\mathcal{C}^{\prime})=\frac{a+b}{{n\choose 2}}$
@@ 5 Experimental results
In our synthetic experiments, we perform hyperparameter optimization of both LINSCAN and OPTICS (for comparison) on 10 synthetic datasets using a random search of the feature space for 500 trials, applying our spectral filtering to both LINSCAN and OPTICS. We then report the test accuracy of both algorithms on 40 synthetic datasets. The results are as follows:
Algorithm OPTICS LINSCAN
Validation ARI 46.73% 61.48%
Testing ARI 46.40% 64.19%
In particular, even though the parameter space for LINSCAN is much larger than OPTICS (optimizing $\mathrm{minPts}$, $\mathrm{eccPts}$, $\xi$, and $\tau$ compared to just $\mathrm{minPts}$, $\epsilon$, and $\tau$), LINSCAN performed better on both the validation data and the testing data and generalized as well as or better than OPTICS. A sample of the performance of LINSCAN and OPTICS on generated data is given in Figure 5(a).
@@
6 Code availability
Name of Repository: LINSCAN_Public
Contact: adennehy@uchicago.edu
Hardware requirements: CPU
Program language: Python, C
Software required: Python, CPython
Program size: 30 MB
Author Remark: Running the code requires compiling a C function for use in the python script. Instructions are included in the repository.
The source codes are available for downloading at the link:
https://github.com/aj111000/LINSCAN_Public
@@ Acknowledgements
Appendix A Approximation of $KL(P|Q)$
@@
$\log\left|A\right|=\mathrm{tr}(\log(A))$
@@
$=\mathrm{tr}\left(\log(\Sigma_{Q})-\log(\Sigma_{P})\right)$
@@
$\mathrm{tr}\left(\log(AB)\right)$ $=\mathrm{tr}(\log(A))+\mathrm{tr}(\log(B))$
@@
Furthermore, if $\left\|A-I\right\|_{F}<1$, then the sum
@@
converges in $\left\|\cdot\right\|_{F}$ to $\log(A)$. Combining all of this, if $\left\|\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right\|<1$ then
$\mathrm{tr}\left(\log(\Sigma_{Q})-\log(\Sigma_{P})\right)$
$=-\mathrm{tr}\left(\log\left(\Sigma_{Q}^{-1/2}\right)+\log\left(\Sigma_{P}\right)+\log\left(\Sigma_{Q}^{-1/2}\right)\right)$
$=-\mathrm{tr}\left(\log\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}\right)\right)$
$=-\mathrm{tr}\left(\sum_{k=1}^{\infty}(-1)^{k+1}\frac{\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)^{k}}{k}\right)$
$=-\sum_{k=1}^{\infty}(-1)^{k+1}\frac{\mathrm{tr}\left(\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)^{k}\right)}{k}$
$=-\mathrm{tr}\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)+\frac{1}{2}\mathrm{tr}\left(\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)^{2}\right)+o\left(\mathrm{tr}\left(\left(\Sigma_{q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)^{3}\right)\right)$
$=-\mathrm{tr}\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)+\frac{1}{2}\left\|\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right\|_{F}^{2}+o\left(\mathrm{tr}\left(\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)^{3}\right)\right)$
where in the last line we used the fact that $\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I$ is symmetric and for any symmetric matrix $A$
$\mathrm{tr}(A^{2})=\mathrm{tr}(A^{T}A)=\left\|A\right\|_{F}^{2}$
@@
$\mathrm{tr}(\Sigma_{Q}^{-1}\Sigma_{P}-I)=\mathrm{tr}(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I)$
@@
$\frac{1}{2}\log\frac{\left|\Sigma_{Q}\right|}{\left|\Sigma_{P}\right|}+\frac{1}{2}\mathrm{tr}(\Sigma_{Q}^{-1}\Sigma_{P}-I)$
$=\frac{1}{4}\left\|\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right\|_{F}^{2}+o\left(\mathrm{tr}\left(\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)^{3}\right)\right)$
@@
Appendix B Proof of Relaxed Triangle Inequality
@@
$\left(\Sigma_{K}^{-1/2}\Sigma_{Q}^{1/2}-I\right)^{T}\left(\Sigma_{K}^{-1/2}\Sigma_{Q}^{1/2}-I\right)$
@@
$\left\|\Sigma_{K}^{-1/2}\Sigma_{Q}^{1/2}-I\right\|_{2}^{2}$ $={\left\|\left(\Sigma_{K}^{-1/2}\Sigma_{Q}^{1/2}-I\right)^{T}{\Sigma_{K}^{-1/2}\Sigma_{Q}^{1/2}-I}\right\|_{2}}$
@@ Appendix C Algorithms
     if #$\#R_{\epsilon}(x)<\mathrm{minPts}$ then
@@
           Pick $\text{Pick }y\in S$
           if #$\#R_{\epsilon}(y)<\mathrm{minPts}$ then
@@
        if #$\#C_{n}<\mathrm{minPts}$ then
@@
  Input: Neighborhood $N$, core point $p$, queue $S$, $\epsilon>0$, $\mathrm{minPts}\in\mathbb{N}$
@@
     if $d_{\mathrm{min}}(o)=\infty$ (Note this means $o\notin S$then

### v2

@@ 3.3 Behavior of Distance Measure
The proof relies on a significant number of inequalities and is provided in the appendix. The proof proceeds by separating the first two terms of $D(P,K)$ from the last two and showing that each pair individually satisfies the triangle inequality with small additive errors.
Importantly, this shows that for small values of $\epsilon$, $D$ behaves approximately like a metric, which allows us to bound the diameter of any cluster in terms of $\epsilon$ and the number of steps between points in the cluster. This ensures that points whose local neighborhoods are nearly orthogonal are not clustered together. Compare this to the best results proven previously for the approximate triangle inequality of the unmodified KL-Divergence between Gaussians in triangle, which was of exponential order.
@@ 4 Numerical Results
Experiments with synthetic data sets revealed that some clusters identified by LINSCAN may not appear as sufficiently "linear" upon visual inspection (e.g., due to high scatter of data points). Therefore we introduce an additional quality check whereby we compute the covariance matrix of each cluster. In the case of ${\mathbb{R}}^{2}$, we set a minimal threshold $\tau$ on the ratio of the minimum eigenvalue to the maximum eigenvalue of the covariance matrix and remove the groups that do not meet this threshold. For fair comparison, we also apply this filtering step to other clustering algorithms we test.
4.1 Runtime Comparisons
One possible issue with working with a custom distance measure is the cost of calculating all possible distances. In figure 2 we plot the cost of calculating all pairwise distances for datasets of various sizes as $\mathrm{eccPts}$ varies, as well as on a system with a GPU to accelerate the distance computation and one without. We can see that even for large amounts of points, the runtime for calculating both distance measures across all pairs of points is less than a second on average. For comparison, the runtime of the actual clustering algorithm is on the order of 20 seconds on our machine, so although the runtime is higher for LINSCAN’s distance measure, that cost is dwarfed by the core clustering algorithm.
On top of this, if further speedups are required, we can use out-of-the box spatial indexing methods. Using 1, we can lower bound the distance between two distributions by the distance between their means, which means that we can perform an efficient initial step where we filter out pairs whose means are sufficiently far apart before calculating our distance measure on the remaining pairs.
Figure 2: Distance Runtime
(a) Runtime without GPU Acceleration
(b) Runtime with GPU Acceleration
@@ 4.2 Example Datasets
Figure 3 shows an example of synthetic data with two QLCs intersecting at a high angle. Note that unlike DBSCAN, LINSCAN is able to separate the two QLCs. Further, we can see the dependence on point order in ADCN, as the clustering behavior is sensitive to initialization, in contrast to LINSCAN which is fully deterministic and independent of the ordering of points.
Figure 1(b) shows the results of applying LINSCAN to the same data as in Figure 1 and Figure 1(c) shows the results of labeling clusters with spectral ratio greater than $\frac{1}{2}$ as noise (and removing noise points for clarity).
Figure 4 shows the results of applying LINSCAN to real data representing earthquake epicenters in Southern California (fi&ji21). Not only does LINSCAN identify QLCs and remove the "diffuse" background seismicity, but it is also able to identify the clusters at multiple distinct scales by varying $\mathrm{minPts}$. If we try to do the same thing with OPTICS we get multiple clusters, but we fail to form specifically linear clusters. For further examples of the use of LINSCAN on real data, we refer the reader to zou23, in particular figure 1.
We don’t contrast against ADCN on the real data, as getting a representative picture of its performance on a dataset of this size requires applying the algorithm many times from different initializations, due to the dependence of ADCN on the order of the points.
@@
(b) Large Scale LINSCAN with Noise Removed
(c) Small Scale LINSCAN with Noise Removed
(d) Large Scale OPTICS with Noise Removed
(e) Small Scale OPTICS with Noise Removed
@@ 4.3 Measuring Performance
To quantitatively evaluate the algorithm performance, we conducted several tests on synthetic, labeled data. Due to the highly specific nature of the problem we are interested in solving, we are unable to use existing benchmark datasets for clustering algorithms. Each of our synthetic data examples consist of 10 linear clusters, 5 isotropic clusters, and 10 pairs of linear clusters intersecting at angles uniformly distributed in the range $1\pi,.9\pi]$ and separated them from one another in space. An example of one synthetic dataset is given in Figure 5(a).
To score the performance, we use the Adjusted Rand Index from Hubert1985. We include a definition of this in the appendix.
@@ 5 Experimental results
In our synthetic experiments, we perform hyperparameter optimization of both LINSCAN and OPTICS (for comparison) on 10 synthetic datasets using a Tree-structured Parzen Estimator (NIPS2011_86e8f7ab) for 500 trials, applying our spectral filtering to both LINSCAN and OPTICS. We then report the test accuracy of both algorithms on 40 synthetic datasets. We report the means and 95% confidence intervals for both the validation and test data in 5.
Synthetic Experiment Results Algorithm OPTICS LINSCAN Validation ARI $0.34\pm 0.09$ $0.60\pm 0.15$ Testing ARI $0.37\pm 0.20$ $0.51\pm 0.14$
In particular, even though the parameter space for LINSCAN is much larger than OPTICS (optimizing $\mathrm{minPts}$, $\mathrm{eccPts}$, $\xi$, and $\tau$ compared to just $\mathrm{minPts}$, $\epsilon$, and $\tau$), LINSCAN performed better on both the validation data and the testing data and generalized as well as or better than OPTICS. A sample of the performance of LINSCAN and OPTICS on generated data is given in Figure 5. For ease of visualization, the noise points are filtered out.
We also remark that although many clusters which are far in space have similar colors, this is a product of the choice of colormap for plotting. All of the final clusters that both LINSCAN and OPTICS return are spatially connected.
Remark 5.1 (Best practice for choosing hyperparameters).
Here we describe informally the observed behavior of LINSCAN when the more unintuitive parameters are varied, for the benefit of would-be practitioners.
  1. a.
    $\xi$: We observed that varying $\xi$ from the default value of $0.05$ from the original OPTICS paper was almost never beneficial to the final performance.
  2. b.
    $\epsilon$: While theoretically there is a benefit to choosing $\epsilon$ small to reduce the computational overhead of the distance function, the computational cost of the distance function was dwarfed by the cost of the clustering in practice. Thus, it is safe to disregard this parameter (or equivalently set it equal to $\infty$) outside of the largest of datasets.
  3. c.
    $\mathrm{eccPts}$: Varying this parameter will adjust the scale at which QLCs are discovered, as in figure 4. For a given dataset of interest, the authors recommend visually tuning this value on a local subset of points to ensure that the extracted clusters correspond to features of the correct scale.
  4. d.
    $\mathrm{minPts}$: Raising the value of this parameter makes clusters less likely, while reducing the number of noise points which appeared in the final clusters. In practice this was often the hardest parameter to tune.
@@
6 Conclusion
We present a method for detecting linear clusters in noisy data. This is done using a novel distance measure, motivated by KL-divergence between small data-driven Gaussian representations of the points, inside of the OPTICS algorithm. We also prove that our distance measure has more regular local behavior than the standard symmetrized KL Divergence. This approach significantly outperforms the DBSCAN family of algorithms that do not have a priori bias towards lineated clusters. Finally, we have shown our approach is shown to be effective in detecting linear slip faults in seismic data and are currently exploring additional applications of our algorithm in other domains.
@@ Acknowledgements
Appendix A Expansion of KL-Divergence Between Gaussians
@@
$\log\left|A\right|=\operatorname{tr}(\log(A))$
@@
$=\operatorname{tr}\left(\log(\Sigma_{Q})-\log(\Sigma_{P})\right)$
@@
$\operatorname{tr}\left(\log(AB)\right)$ $=\operatorname{tr}(\log(A))+\operatorname{tr}(\log(B))$
@@
Furthermore, if ${\left\|A-I\right\|}<1$ for a submultiplicative norm $\left\|\cdot\right\|$}, then the sum
@@
converges to $\log(A)$. Combining all of this, if ${\left\|\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right\|}<1$ then
$\operatorname{tr}\left(\log(\Sigma_{Q})-\log(\Sigma_{P})\right)$
$=-\operatorname{tr}\left(\log\left(\Sigma_{Q}^{-1/2}\right)+\log\left(\Sigma_{P}\right)+\log\left(\Sigma_{Q}^{-1/2}\right)\right)$
$=-\operatorname{tr}\left(\log\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}\right)\right)$
$=-\operatorname{tr}\left(\sum_{k=1}^{\infty}(-1)^{k+1}\frac{\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)^{k}}{k}\right)$
$=-\sum_{k=1}^{\infty}(-1)^{k+1}\frac{\operatorname{tr}\left(\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)^{k}\right)}{k}$
$=-\operatorname{tr}\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)+\frac{1}{2}\operatorname{tr}\left(\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)^{2}\right)+o\left(\operatorname{tr}\left(\left(\Sigma_{q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)^{3}\right)\right)$
$=-\operatorname{tr}\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)+\frac{1}{2}{\left\|\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right\|}_{F}^{2}+o\left(\operatorname{tr}\left(\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)^{3}\right)\right)$
where in the last line we used the fact that $\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I$ is symmetric and for any symmetric matrix AA
$\operatorname{tr}(A^{2})=\operatorname{tr}(A^{T}A)={\left\|A\right\|}_{F}^{2}$
@@
$\operatorname{tr}(\Sigma_{Q}^{-1}\Sigma_{P}-I)=\operatorname{tr}(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I)$
@@
$\frac{1}{2}\log\frac{\left|\Sigma_{Q}\right|}{\left|\Sigma_{P}\right|}+\frac{1}{2}\operatorname{tr}(\Sigma_{Q}^{-1}\Sigma_{P}-I)$
$=\frac{1}{4}{\left\|\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right\|}_{F}^{2}+o\left(\operatorname{tr}\left(\left(\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}-I\right)^{3}\right)\right)$
@@
Appendix B Proof of Lemma 3.1
Lemma.
Let $P={\mathcal{N}}(\mu_{P},\Sigma_{P})$ and $Q={\mathcal{N}}(\mu_{Q},\Sigma_{Q})$ be Gaussians. Then,
  1. a.
    $D$ is symmetric, meaning
    $D(P,Q)=D(Q,P)$
  2. b.
    $D(P,Q)=0$ iff $P=Q$ (in particular $D(P,P)=0$)
Proof.
Proof:
  1. a.
    Trivial
  2. b.
    Note that by the definition of $D$,
    $D(P,Q)=0$ $\iff{\left\|\mu_{P}-\mu_{Q}\right\|}_{\Sigma_{Q}^{-1}}={\left\|\mu_{P}-\mu_{Q}\right\|}_{\Sigma_{P}^{-1}}=0\text{ and }\Sigma_{Q}^{-1/2}\Sigma_{P}\Sigma_{Q}^{-1/2}=\Sigma_{P}^{-1/2}\Sigma_{Q}\Sigma_{P}^{-1/2}=I$
    Since we assume all of the covariance matrices are invertible, the first equalities hold iff $\mu_{P}=\mu_{Q}$. Similarly, the second equalities hold iff $\Sigma_{P}=\Sigma_{Q}$. Hence, $D(P,Q)=0$ iff $P=Q$
Appendix C Proof of Relaxed Triangle Inequality
Theorem.
Let $\epsilon>0$. If $D(P,Q),D(Q,K)\leq\epsilon$, then
@@ Appendix C Proof of Relaxed Triangle Inequality
$(\Sigma_{K}^{-1/2}\Sigma_{Q}^{1/2}-I)^{T}(\Sigma_{K}^{-1/2}\Sigma_{Q}^{1/2}-I)$
@@
${\left\|\Sigma_{K}^{-1/2}\Sigma_{Q}^{1/2}-I\right\|}_{2}^{2}$ $={{\left\|(\Sigma_{K}^{-1/2}\Sigma_{Q}^{1/2}-I)^{T}(\Sigma_{K}^{-1/2}\Sigma_{Q}^{1/2}-I)\right\|}_{2}}$
@@ Appendix D Algorithms
  if $R_{\epsilon}(x)<\mathrm{minPts}$ then
@@
    $\text{Pick }y\in S$
    if $R_{\epsilon}(y)<\mathrm{minPts}$ then
@@
   if $C_{n}<\mathrm{minPts}$ then
@@
Input: Neighborhood NN, core point $p$, queue $S$, $\epsilon>0$, $\mathrm{minPts}\in{\mathbb{N}}$
@@
  if $d_{\mathrm{min}}(o)=\infty$ (Note this means $o\notin S$) then
@@
Appendix E Rand Index and Adjusted Rand Index
Definition E.1 (Rand Index).
Let $X=\left\{x_{1},...,x_{n}\right\}$ and consider two partitions $\mathcal{C}=\left\{C_{1},...,C_{m}\right\}$ and $\mathcal{C}^{\prime}=\left\{C^{\prime}_{1},...,C^{\prime}_{n}\right\}$ of $X$, i.e. $C_{i}\subseteq X$ and $C_{i}^{\prime}\subseteq X$ for all $i$ and
$X=\bigcup_{i=1}^{m}C_{i}=\bigcup_{i=1}^{n}C^{\prime}_{i}$
with
$C_{i}\cap C_{j}=C_{i}^{\prime}\cap C_{j}^{\prime}=\emptyset$
for all $i\neq j$. Let
$a:=\#\left\{(x,y)\in X\times X:x\neq y,\exists i,j\text{ s.t. }x,y\in C_{i},x,y\in C_{j}^{\prime}\right\}$
and
$b:=\#\left\{(x,y)\in X\times X:x\neq y,\exists i,j,k,l\text{ s.t. }i\neq j,k\neq l,x\in C_{i},x\in C_{k}^{\prime},y\in C_{j},y\in C_{l}^{\prime}\right\}$
$a$ is the number of pairs of elements of $X$ such that both elements are in the same cluster in $\mathcal{C}$ and $\mathcal{C}^{\prime}$ and $b$ is the number of pairs of elements of $X$ such that both elements are in different clusters in both $\mathcal{C}$ and $\mathcal{C}^{\prime}$. Then, the Rand Index is given by
$R(X,\mathcal{C},\mathcal{C}^{\prime})=\frac{a+b}{\binom{n}{2}}$
Code availability
The source codes are available for downloading at the link:
https://github.com/aj111000/LINSCAN_Public
We produced the code and images in this work using NumPy, JAX, Matplotlib, scikit-learn, and Optuna (NumPy, jax, matplotlib, sklearn_api, optuna).